Arrange the following compounds in order of increa
Number 26 only
Answer #1

KHP is acid and NaOH is base.
We have acid base reaction.

KHP + NaOH ---> H2O + K+(aq) + Na+ (aq) + P2-(aq)

so KHP and NaoH react in 1;1 ratio.

KH moles = mass of KHP / molar mass of KHP = 0.121 /204.2 = 0.00059255

Now NaoH moles = KHP moles = 0.00059255

Mollarity fo NaOH = moles of NaoH / volume of NaoH in L

M of NaoH = ( 0.00059255) / ( 0.0325) = 0.01823 M

thus concentrationof NaoH= 0.01823 M